# PlanGenLLMs: A Modern Survey of LLM Planning Capabilities

### v2

@@ 2 LLM Planning Foundations (Tables 1-3)
Commonly used pretrained LLMs (e.g., GPT-4 Achiam et al. (2023), Claude 3 Anthropic (2024), Llama 2 Touvron et al. (2023)) are not specifically trained for planning tasks, and prompt-based methods, which do not update model parameters, cannot fundamentally improve performance in these areas Chen et al. (2023a); Wang et al. (2024). Fine-tuning, either focused on planning-specific tasks or broader agentic capabilities, enhances planning correctness by directly updating LLM parameters.
@@ 3 Criterion I: Completeness (Table 4)
A plan is correct if it achieves the goal within a fixed budget while avoiding excessive complexity and infinite loops (e.g., finishing booking a flight ticket to San Francisco within five hours meets the criteria for correctness). To ensure correctness, the LLM must work with classical sound and complete solvers Guan et al. (2023); Hao et al. (2024a). Also, the LLM has to accurately translate the domain and problem into the specific format (e.g., PDDL), required by these solvers Guan et al. (2023).
@@ 4 Criterion II: Executability (Tables 5-6)
Object Grounding Object grounding ensures the LLM planner uses objects available in the current environment when generating plans. E.g., if a stove is present but a microwave is not, the planner should correctly select the stove to heat a pancake and exclude the unavailable microwave from its plan. The simplest way to do this is by feeding observed or available objects into the planner via prompts (Huang et al., 2022b; Song et al., 2023; Lin et al., 2023a; Singh et al., 2023) or neural embeddings Sharma et al. (2021); Ahn et al. (2022). In partially observed environments, where some object information are uncertain (e.g., needing to clean a cup that could be in a cabinet, drawer, or fridge), the planner can generate multiple possible plans, one for each scenario, and select the first feasible one (Prasad et al., 2023; Dagan et al., 2023; Zhao et al., 2024). Sun et al. (2024) takes a different approach, first generating a plan with placeholders for objects, then filling in the blanks with observed objects during execution.
Action Grounding Action grounding ensures all actions in a plan can actually be executed in the current environment. E.g., the planner should not include an action like “draw a star” if the robot arm cannot perform such a complex operation. Instead, the task should be decomposed into simpler supported actions, such as multiple steps of “draw a line”. Like object grounding, the simplest way is to explicitly list all admissible actions in LLMs’ inputs Singh et al. (2023). If a step goes beyond the executor’s capabilities (e.g., combining multiple allowed actions into one), the LLM planner should be reprompted to break it down until every step is executable (Prasad et al., 2023).
@@ 5 Criterion III: Optimality (Table 7)
LLM + Optimizer   It combines the LLM, which turns user requests into symbolic optimization problems, with an optimizer that solves them and finds the best solution Ju et al. (2024); Hao et al. (2024b). For example, TTG (Ju et al., 2024) uses the LLM to convert travel planning requests of minimum total costs into Mixed Integer Linear Programming problems, then runs an optimizer such as SCIP (Bestuzheva et al., 2021) to provide the optimal plan. Compared to LLM + classical planners, where LLMs define the domain and problem in a formal representation (e.g., PDDL), LLM + optimizers ensure optimal solutions by leveraging LLMs to formulate constrained optimization problems and classical optimizers to solve them. Unlike classical planners, which typically rely on search algorithms, heuristics and logical deductions and may not guarantee optimality Russell and Norvig (2016), optimizers, often using gradient-based methods (e.g., Newton’s methods), can guarantee optimal solutions Boyd and Vandenberghe (2004).
A* Search-Based Methods A* search finds the lowest-cost optimal solution, particularly when using admissible heuristics that do not overestimate the actual cost to the goal. This makes it a natural choice for LLM-based planners to achieve optimality. ToolChain* Zhuang et al. (2023) combines A* tree search with an LLM, which suggests next steps and estimates heuristic scores, to create plans with the fewest tool API calls. SayCanPay (Hazra et al., 2024) uses A* search with LLMs to generate the shortest possible plans. Beyond A* (Lehnert et al., 2024) trains a Transformer model, Searchformer, to mimic A* search paths for complex tasks like Maze navigation and Sokoban puzzles, optimizing for the fewest steps. Besides A* search, other search algorithms (e.g., DFS and MCTS) could also be used to find optimal solutions, although without guarantee.
@@ 8 Criterion VI: Efficiency (Table 11)
To decrease LLM and world model calls, several tricks are used: (1) generating the entire plan in one shot instead of step-by-step to reduce redundant prompts Hu et al. (2023b); Sun et al. (2024); Gonzalez-Pumariega et al. (2024); (2) checking feasibility by world models only at the end of each subgoal, not after every action (Sun et al., 2024; Gonzalez-Pumariega et al., 2024); (3) merging plans with the same prefix actions or subgoals to avoid duplicate world model checks when sampling multiple plans (Hu et al., 2023b); and (4) in tree search-based methods, querying the LLM once to generate a successor function and a goal check function. The successor function, which can produce all possible actions and states, provides state transitions based on the current state and selected action. The goal-check function determines whether the current state is the desired final goal. This approach avoids repeated LLM and world model calls at each node (Katz et al., 2024).

### v3

@@ 2 LLM Planning Foundations (Tables 1-3)
Current LLMs are not specifically trained for agentic tasks like planning, and prompt-based methods, which do not update model parameters, cannot fundamentally improve performance in these areas Chen et al. (2023a); Wang et al. (2024). Fine-tuning, either focused on planning-specific tasks or broader agentic capabilities, enhances planning correctness by directly updating LLM parameters.
@@ 3 Criterion I: Completeness (Table 4)
A plan is correct if it achieves the goal within a fixed budget while avoiding excessive complexity and infinite loops. To ensure correctness, the LLM must work with classical sound and complete solvers Guan et al. (2023); Hao et al. (2024a). Also, the LLM has to accurately translate the domain and problem into the specific format (e.g., PDDL), required by these solvers Guan et al. (2023).
@@ 4 Criterion II: Executability (Tables 5-6)
Object Grounding Object grounding ensures the LLM planner uses objects available in the current environment when generating plans. For example, if a stove is available but a microwave is not, the planner should choose the stove to heat a pancake and disregard the unavailable microwave in its plan. The simplest way to do this is by feeding observed or available objects into the planner via prompts (Huang et al., 2022b; Song et al., 2023; Lin et al., 2023a; Singh et al., 2023) or neural embeddings Sharma et al. (2021); Ahn et al. (2022). In partially observed environments, where some object information are uncertain (e.g., needing to clean a cup that could be in a cabinet, drawer, or fridge), the planner can generate multiple possible plans, one for each scenario, and select the first feasible one (Prasad et al., 2023; Dagan et al., 2023; Zhao et al., 2024). Sun et al. (2024) takes a different approach, first generating a plan with placeholders for objects, then filling in the blanks with observed objects during execution.
Action Grounding Action grounding ensures all actions in a plan can actually be executed in the current environment. E.g., the planner should avoid including actions such as ‘draw a star’ if the robot arm cannot perform such a complex operation. Instead, the task should be decomposed into simpler, supported actions, such as multiple steps of ’draw a line’. Like object grounding, the simplest way is to explicitly list all admissible actions in LLMs’ inputs Singh et al. (2023). If a step goes beyond the executor’s capabilities (e.g., combining multiple allowed actions into one), the LLM planner should be reprompted to break it down until every step is executable (Prasad et al., 2023).
@@ 5 Criterion III: Optimality (Table 7)
LLM + Optimizer   It combines the LLM, which turns user requests into symbolic optimization problems, with an optimizer that solves them and finds the best solution Ju et al. (2024); Hao et al. (2024b). For example, TTG (Ju et al., 2024) uses the LLM to convert travel planning requests of minimum total costs into Mixed Integer Linear Programming problems, then runs an optimizer such as SCIP (Bestuzheva et al., 2021) to provide the optimal plan. LLM + Optimizers ensure optimal solutions by leveraging LLMs to formulate constrained optimization problems and classical optimizers to solve them. In contrast to classical planners, which typically rely on search algorithms, heuristics, and logical deduction and may not guarantee optimality Russell and Norvig (2016), optimizers, often based on gradient-based methods (e.g., Newton’s methods), can guarantee optimal solutions Boyd and Vandenberghe (2004).
A* Search-Based Methods  A* search finds the lowest-cost optimal solution, particularly when using admissible heuristics that do not overestimate the actual cost to the goal. This makes it a natural choice for LLM-based planners to achieve optimality. ToolChain* Zhuang et al. (2023) combines A* tree search with an LLM, which suggests next steps and estimates heuristic scores, to create plans with the fewest tool API calls. SayCanPay (Hazra et al., 2024) uses A* search with LLMs to generate the shortest possible plans. Beyond A* (Lehnert et al., 2024) trains a Transformer model, Searchformer, to mimic A* search paths for complex tasks like Maze navigation and Sokoban puzzles, optimizing for the fewest steps. Besides A* search, other search algorithms (e.g., DFS and MCTS) could also be used to find optimal solutions, although without guarantees.
@@ 8 Criterion VI: Efficiency (Table 11)
To reduce the number of LLM and world model calls, several tricks are used: (1) generating the entire plan in one shot instead of step-by-step to reduce redundant prompts Hu et al. (2023b); Sun et al. (2024); Gonzalez-Pumariega et al. (2024); (2) checking feasibility by world models only at the end of each subgoal, not after every action (Sun et al., 2024; Gonzalez-Pumariega et al., 2024); (3) merging plans with the same prefix actions or subgoals to avoid duplicate world model checks when sampling multiple plans (Hu et al., 2023b); and (4) in tree search-based methods, querying the LLM once to generate a successor function and a goal-check function. The successor function lists all possible actions and resulting states based on the current state, while the goal-check function determines if a state meets the goal. This approach avoids repeated LLM and world model calls at each node (Katz et al., 2024).